OpenStudy (anonymous):

A cylindrical vase has a diameter of 6in. At the bottom of the vase, there are 6 marbles, each of diameter of 1in. the vase is filled with water up to a height of 12in. What is the volume of water in the vase. A. 102pi in^3 B. 106pi in^3 C. 107pi in^3 D. 108pi in^3

OpenStudy (anonymous):

total volume - volume of marbles = volume of water Total volume =12 (6/2)^2 pi volume of marbles = 6(1^2 pi ) volume of water = total - marbles =12 (6/2)^2 pi -6(1^2 pi ) =102 pi

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

Volume marbles 6*4/3*(1/2)^3*pi = pi

OpenStudy (anonymous):

So the answer is C

OpenStudy (anonymous):

yes yes yes Sorry for making mistakes